# What-if Analysis for Business Users: Current Practices and Future Opportunities

### v3

@@ Abstract.
What-if analysis (WIA), crucial for making data-driven decisions, enables users to understand how changes in variables impact outcomes and explore alternative scenarios. However, existing WIA research focuses on supporting the workflows of data scientists or analysts, largely overlooking significant non-technical users, like business users. We conduct a two-part user study with 22 business users (marketing, sales, product, and operations managers). The first study examines existing WIA techniques employed, tools used, and challenges faced. Findings reveal that business users perform many WIA techniques independently using rudimentary tools due to various constraints. We implement representative WIA techniques identified previously in a visual analytics prototype to use as a probe to conduct a follow-up study evaluating business users’ practical use of the techniques. These techniques improve decision-making efficiency and confidence while highlighting the need for better support in data preparation, risk assessment, and domain knowledge integration. Finally, we offer design recommendations to enhance future business analytics systems.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Conference on Human Factors in Computing Systems; April 26–May 01, 2025; Yokohama, Japanisbn: 978-1-4503-XXXX-X/18/06submissionid: 2546ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studiesccs: Human-centered computing Visual analyticsccs: Human-centered computing Graphical user interfaces
@@ 1. Introduction
What-if analysis (WIA) is crucial for making data-driven decisions (Gathani et al., 2022; Dimara et al., 2021). For example, business managers may want to understand which marketing activities (e.g., sending emails, making phone calls, giving out free trials, running ad campaigns) should be prioritized to increase customer acquisition rates (Gathani et al., 2022); public health researchers might be interested in the effect of changes in socio-economic factors (e.g., smartphone use is maximized or sleep quality is low) on adolescent obesity rates (Kim et al., 2019); operations analysts may want to identify the optimal time complexities of various sub-algorithms used in a solver (e.g., cut generation, global constraint, local search move) needed to reduce the solver’s overall computation time (Van Cauwelaert et al., 2017). To answer these questions, WIA is essential for identifying key data variables that influence specific outcomes, simulating various scenarios to understand how changes in data variables impact predictions of outcomes, and determining optimal data variable values needed to achieve desired target outcomes. For example, in marketing, given a target outcome of increasing customer acquisition by 5%, WIA can be accomplished by first identifying key variables such as running ad campaigns and making phone calls, building models that predict customer acquisitions based on these variables, and trying alternative scenarios involving different numbers of campaigns and phone calls. Informed decisions can then be made based on the observations on the resulting acquisition rates in these scenarios.
@@
In such examples across many different domains, WIA plays an important role that cannot be fulfilled by other data analysis methods. Similar to many analytic methods, WIA involves building models (e.g., statistical, probabilistic, neural network, optimization, etc.) of independent and dependent variables. However, unlike confirmatory methods such as hypothesis testing, WIA is more exploratory in nature: it involves changing variable values in multiple scenarios and comparing possible outcomes without significance testing. On the other hand, exploratory data analysis (EDA) (Tukey, 1977), which is concerned with understanding trends and patterns in a dataset, is not appropriate for simulating alternative scenarios and predicting their outcomes (Dimara et al., 2021; Oral et al., 2023) to make data-driven decisions. With data growing exponentially (Duarte, 2024), limitations of human memory and cognitive capacity in generating hypotheses, testing scenarios, and predicting outcomes also exacerbate (Gathani et al., 2022). WIA thus addresses an important need that is not supported by methods described, and a growing body of work is dedicated to developing WIA tools for various domains (Wexler et al., 2019; Grafberger et al., 2023; Hohman et al., 2019; Sacha et al., 2018; Gallo et al., 2018; Torsney-Weir et al., 2018; Eric Newburger, 2023).
@@
Despite these efforts to support WIA, existing research mostly focuses on technical users like data analysts and scientists, and overlooks non-technical users who typically lack the expertise or background in coding, statistics, and algorithmic modeling like healthcare experts, urban planners, or law officials (Gathani et al., 2022; Crisan and Correll, 2021; Dimara et al., 2021; Oral et al., 2023). WIA is often practiced by non-technical users because they possess critical domain knowledge that technical users may lack. For example, in WIA, it is often important to brainstorm alternative scenarios and account for potential trade-offs between variables (Dimara et al., 2021), which require extensive domain-specific experiences and intuitions that are not found in professional data analysts. Additionally, non-technical users may also build upon and extend data-driven insights generated by technical users towards specific desired outcomes (Crisan and Correll, 2021; Bhattacharya et al., 2024). Understanding their current WIA practices and pain points is essential for developing better tools that effectively integrate domain knowledge with WIA techniques.
@@
In this paper, we choose to focus on a specific key group of non-technical users conducting WIA: business users [^1] In the remainder of the paper, we use the terms business users and users interchangeably.—such as sales managers, marketing managers, product managers, and operations managers. These users represent a significantly large and influential demographic group (Report, 2022b, a) that regularly makes data-driven decisions without formal training in technical fields like computer science, machine learning, or statistics. They are also the primary audience for self-service business intelligence (BI) tools which constitute a multi-billion dollar industry (Insights, 2024) and stand as the most prominent commercial application of visual data analytics (Software, 2024). Since little is known about how business users employ WIA and the challenges they encounter, we seek to answer the following research questions:
@@
RQ1: What techniques and tools do business users employ to perform WIA for making data-driven decisions?
@@
RQ2: How do they perceive and interact with advanced WIA techniques, and what future opportunities exist for improving these techniques?
@@
We answer these by conducting a two-part user study with 22 business users. For RQ1, we conduct a semi-structured interview study to understand their WIA workflows and challenges when making data-driven decisions. In addition to discussing their own experiences, we ask participants to also walk us through a common business use case of maximizing sales by making spending decisions in various advertising channels, often referred to as marketing mix modeling. Our findings reveal that business users employ various WIA techniques, but rely on rudimentary methods like spreadsheets which are insufficient for making data-driven decisions. Compared to prior studies on data workers’ generic analysis practices (Dimara et al., 2021, 2021; Tory et al., 2021), our study uncovers a specific set of WIA techniques adopted by business users (e.g., sensitivity analysis, driver importance analysis, segmentation analysis) and the associated challenges they face. Moreover, business users prefer performing WIA independently, without relying on data analysts, due to (1) the limited availability of data analysts in enterprises, (2) the inefficiency of communicating with data analysts, (3) the business pressure to make quick decisions, and (4) the importance of incorporating domain knowledge, which data analysts often lack. These findings suggest that currently there is inadequate support for business users to perform WIA making data-informed decisions.
@@
The first interview study informs us of various WIA techniques employed by business users and the challenges they encounter with them. We address RQ2 by first implementing four representative WIA techniques identified in the interview study (i.e., driver importance analysis, sensitivity analysis, goal-seeking analysis, and constrained analysis) in a visual analytics prototype. We then conduct a follow-up study with the same participants, asking them to use this prototype as a probe to make decisions for the same marketing mix modeling use case as in the interview study. Through the study, we identified the potential benefits and opportunities of WIA techniques for business users. Our findings show that they improve business users’ decision-making speed and confidence, highlighting their strong interest in adopting additional advanced analytics techniques for informing their decisions. However, participants also expressed opportunities for more support in preparing data (e.g., consolidating data from multiple sources, consistently defining KPI goals), assessing prediction risks (e.g., understanding confidence and trust in the predictions), and incorporating domain knowledge (e.g., capturing limited budgets, volatile ecological and market conditions such as COVID, war, and inflation).
@@
  • An interview study identifying WIA techniques, tools, and challenges that business users face when making data-driven decisions.
  • A task-based study gathering business users’ hands-on experience and feedback on the benefits and opportunities of using and enhancing WIA techniques for their use in making data-driven decisions.
  • Design recommendations for future business analytics systems to better support business users in advanced analytics.
@@ 2. Related Work
Our work relates to prior research on other users performing data analysis, use of enterprise dashboards, advanced analytics tools, and organizational decision-making.
2.1. Understanding Users Performing Data Analysis
Prior work has examined how different types of users perform data analysis. We categorize them along two critical dimensions: task (data understanding and sense-making vs. improving outcomes and decision-making) and technical expertise (having formal technical training vs. no formal training), as depicted in Figure 1.
Figure 1. Layout of data-analyzing users studied in the literature. We map the user groups considering the dimensions of task (x-axis) and technical expertise (y-axis). Our research focuses on business users, who are domain experts lacking technical (e.g., programming, statistics, data science, etc.) background but using data analysis to make better business decisions.
@@ 2.1. Understanding Users Performing Data Analysis
Most studies focus on users with technical expertise (quadrant I), like data analysts, data scientists, and business analysts, who mainly support decisions rather than make them. For example, Kandel et al. (Kandel et al., 2012) identify challenges in analysts’ data pipelines, while Alspaugh et al. (Alspaugh et al., 2019) explore challenges in exploratory analysis. Demiralp et al. (Demiralp et al., 2017) study data scientists’ exploratory analysis, and Zhang et al. investigate business analysts’ needs in generating reports (Zhang et al., 2020) and narratives (Zhang et al., 2022). Although we find business users face similar challenges (e.g., the need to bridge the programming gap for analysts who are less savvy in programming, and the challenges of data integration), our findings highlight their need for data-driven WIA to explore alternative scenarios and the unique challenges related to making decisions in terms of time pressure, uncertainty, and domain knowledge. Newburger and Elmquist study statisticians’ decision-making with a focus on visualization (Eric Newburger, 2023). While their work shares some similarities with ours, we delve deeper into the mechanisms, tools, and challenges of their advanced analyses.
@@
Bartram et al. (Bartram et al., 2021) study “data workers” (quadrants III and IV) with less formal expertise who perform data analysis as part of their daily tasks. Previous work often focuses on data workers in non-business domains, such as political science (South et al., 2020), community leadership (Jasim et al., 2021), and law enforcement (Kang and Stasko, 2012), but their tasks are more about data exploration and sense-making (Klein et al., 2006; Russell et al., 1993; Klein et al., 2007; Pirolli and Card, 2005) than making decisions involving simulating alternative scenarios and action selection (Lunenburg, 2010; Bruch and Feinberg, 2017; Chen, 2008; Oral et al., 2023). Current research also often prioritizes unstructured data, like text. For example, Edge et al. (Edge et al., 2018) report on bringing AI within unstructured sources like social media, news, and cyber intelligence. However, a large number of use cases primarily involve structured data tables, which are underrepresented in studies. Bhattacharya et al. (Bhattacharya et al., 2024) comes closest in empowering non-technical domain experts, particularly healthcare professionals, by enhancing AI models through diverse model explanations (local and global). However, their focus is on improving models and predictions rather than supporting decision-making processes. On contrary, this paper examines how business users explore data scenarios using WIA techniques to practically make decisions.
@@ 2.2. Use of Dashboards in Enterprises
Previous work investigates the role of visualization and dashboards in enterprise data analysis and decision-making. For example, Newburger and Elmqvist (Eric Newburger, 2023) explore visualization for statisticians, while studies find that business users, like sales managers, frequently use dashboards (Velcu-Laitinen and Yigitbasioglu, 2012; Bera, 2016; Noonpakdee et al., 2018). However, dashboards are mainly used for communicating findings to stakeholders, not for making data-driven decisions. Dimara and Stasko (Dimara and Stasko, 2021) note the lack of studies on decision-making in visualization research, and other works (Dimara et al., 2021) show how visualizations can be integrated into long-term decisions. Yet, business users face challenges in creating dashboards and conducting analysis necessary for making decisions (Sarikaya et al., 2018; Tory et al., 2021) due to incomplete information shown in dashboard and limited interactivity, which prevents users from drilling into the data to explore different scenarios or conduct simulations. Therefore, visualization alone without data-driven analytics is insufficient, leading us to focus on how advanced analytics can support or hinder making decisions.

### v4

# What-if Analysis for Business Professionals: Current Practices and Future Opportunities

@@ Abstract.
What-if analysis (WIA) is essential for data-driven decision-making, allowing users to assess how changes in variables impact outcomes and explore alternative scenarios. Existing WIA research primarily supports the workflows of data scientists and analysts, and largely overlooks business professionals who engage in WIA through non-technical means. To bridge this gap, we conduct a two-part user study with 22 business professionals across marketing, sales, product, and operations roles. The first study examines their existing WIA practices, tools, and challenges. Findings reveal that business professionals perform many WIA techniques independently using rudimentary tools due to various constraints. We then implement representative WIA techniques in a visual analytics prototype and use it as a probe to conduct a follow-up study evaluating business professionals’ practical use of the techniques. Results show that these techniques improve decision-making efficiency and confidence while underscoring the need for better support in data preparation, risk assessment, and domain knowledge integration. Finally, we offer design recommendations to enhance future business analytics systems.
@@
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713672isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studiesccs: Human-centered computing Visual analyticsccs: Human-centered computing Graphical user interfaces
@@ 1. Introduction
What-if analysis (WIA) is crucial for making data-driven decisions (Gathani et al., 2022; Dimara et al., 2021). For example, business managers may want to understand which marketing activities (e.g., sending emails, making phone calls, giving out free trials, running ad campaigns) should be prioritized to increase customer acquisition rates (Gathani et al., 2022); public health researchers might be interested in examining the effects of socio-economic factors (e.g., maximized smartphone use or low sleep quality) on adolescent obesity rates (Kim et al., 2019); operations analysts may want to identify the optimal time complexities of various sub-algorithms used in a solver (e.g., cut generation, global constraint, local search) to reduce the solver’s overall computation time (Van Cauwelaert et al., 2017). To answer these questions, WIA is essential for identifying key data variables that influence specific outcomes, simulating various scenarios to understand how changes in data variables impact outcome predictions, and determining optimal data variable values to achieve desired target outcomes. For example, in marketing, WIA can be used to achieve a target outcome: increasing customer acquisition by 5%. The process begins by first identifying key variables such as ad campaigns and phone calls, and building models that predict customer acquisitions based on these variables. By exploring alternative scenarios with varying numbers numbers of campaigns and phone calls, businesses can assess their impact on acquisition rates. These insights enable data-driven decision-making to optimize marketing strategies effectively.
@@
In such examples across many different domains, WIA plays an important role that traditional data analysis methods cannot fulfill. Similar to many analytics methods, WIA involves building models (e.g., statistical, probabilistic, neural network, optimization) that represents relationships between independent and dependent variables. However, unlike confirmatory methods such as hypothesis testing, WIA is inherently exploratory: it involves changing variable values in multiple scenarios and comparing possible outcomes without significance testing. On the other hand, exploratory data analysis (EDA) (Tukey, 1977), which focuses on understanding trends and patterns in a dataset, is not appropriate for simulating alternative scenarios and predicting their outcomes to make data-driven decisions (Dimara et al., 2021; Oral et al., 2023). As data continues to grow exponentially (Duarte, 2024), limitations of human memory and cognitive capacity in generating hypotheses, testing scenarios, and predicting outcomes also exacerbate (Gathani et al., 2022). WIA thus addresses a critical need that is not supported by these methods, and a growing body of work is dedicated to developing WIA tools for various domains (Wexler et al., 2019; Grafberger et al., 2023; Hohman et al., 2019; Sacha et al., 2018; Gallo et al., 2018; Torsney-Weir et al., 2018; Eric Newburger, 2023).
@@
Despite efforts to support WIA, existing research mostly focuses on data analysts and scientists with technical expertise in coding, statistics, and algorithmic modeling. It overlooks domain professionals like healthcare experts, urban planners, or law officials (Gathani et al., 2022; Crisan and Correll, 2021; Dimara et al., 2021; Oral et al., 2023) who rely on no-code methods like graphical user interfaces and direct manipulation. These non-technical domain professionals frequently engage in WIA because they possess critical domain knowledge. For example, brainstorming alternative scenarios and evaluating potential trade-offs between variables (Dimara et al., 2021) require extensive domain-specific experience and intuition—skills that professional data analysts may lack. Additionally, domain professionals play a key role in translating data-driven insights generated by data analysts and scientists into actionable strategies that align with their goals (Crisan and Correll, 2021; Bhattacharya et al., 2024). Understanding their current practices and pain points is essential for developing better tools that effectively integrate domain knowledge with WIA techniques.
@@
In this paper, we focus on a key group of domain professionals conducting WIA: business professionals  [^1] In the remainder of the paper, we use the terms business professionals and professionals interchangeably.—such as sales managers, marketing managers, product managers, and operations managers. These professionals represent a significantly large and influential demographic group (Report, 2022b, a) that regularly makes data-driven decisions without employing technical methods like coding and often has limited formal training in technical fields like computer science, machine learning, or statistics. They are also the primary audience for self-service business intelligence (BI) tools, which constitute a multi-billion dollar industry (Insights, 2024) and one of the most prominent commercial applications of visual data analytics (Software, 2024). Since little is known about how business professionals employ WIA and the challenges they face, we seek to answer the following research questions:
@@
RQ1: What techniques and tools do business professionals employ to perform WIA for making data-driven decisions?
@@
RQ2: How do they perceive and interact with advanced WIA techniques, and what future opportunities exist for improving them?
@@
We answer these by conducting a two-part user study with 22 business professionals. For RQ1, we conduct a semi-structured interview study to understand their WIA workflows and challenges when making data-driven decisions. In addition to discussing their own experiences, we ask participants to walk us through a common business use case: maximizing sales by making spending decisions in various advertising channels, often referred to as marketing mix modeling. Our findings reveal that business professionals employ various WIA techniques but rely on rudimentary methods such as those provided by spreadsheet applications, which are insufficient for making data-driven decisions. Compared to prior studies on data workers’ generic analysis practices (e.g., (Dimara et al., 2021; Tory et al., 2023)), our study uncovers a specific set of WIA techniques adopted by business professionals (e.g., sensitivity analysis, driver importance analysis, segmentation analysis) and the associated challenges they face. Moreover, business professionals prefer performing WIA independently, without relying on data analysts, due to (1) the limited availability of data analysts in enterprises, (2) the inefficiency of communicating with data analysts, (3) the business pressure to make quick decisions, and (4) the importance of incorporating domain knowledge, which data analysts often lack. These findings suggest that there is inadequate support for business professionals to perform WIA for making data-informed decisions.
@@
The first interview study informs us of various WIA techniques employed by business professionals and the challenges they encounter with them. We address RQ2 by first implementing four representative WIA techniques identified from the initial interview study (i.e., driver importance analysis, sensitivity analysis, goal-seeking analysis, and constrained analysis) in a visual analytics prototype. We then conduct a follow-up study with the same participants, asking them to use this prototype as a probe to make decisions for the same marketing mix modeling use case as in the interview study. Through the study, we identified the potential benefits and opportunities of WIA techniques for business professionals. Our findings show that they improve business professionals’ decision-making speed and confidence, highlighting their strong interest in adopting additional advanced analytics techniques for informing their decisions. However, participants also expressed opportunities for more support in preparing data (e.g., consolidating data from multiple sources, consistently defining KPI goals), assessing prediction risks (e.g., understanding confidence and trust in predictions), and incorporating domain knowledge (e.g., capturing limited budgets, accounting for volatile ecological and market conditions such as COVID and inflation).
@@
  • An interview study identifying WIA techniques, tools, and challenges that business professionals face when making data-driven decisions.
  • A task-based study gathering business professionals’ hands-on experience and feedback on the benefits and opportunities of using and enhancing WIA techniques for their use in making data-driven decisions.
  • Design recommendations for future business analytics systems to better support business professionals in advanced analytics.
@@ 2. Related Work
Our work relates to prior research on other professionals performing data analysis, use of enterprise dashboards, advanced analytics tools, and organizational decision-making.
2.1. Understanding Professionals Performing Data Analysis
Prior work has examined how different types of professionals perform data analysis. Table 1 summarizes the data, high-level tasks, and levels of technical engagement (i.e., low, medium, high) in some related works examining data analysis across various audience roles, providing context for the landscape.
Paper
Job Role Data High-Level Task Technical Engagement
Kandel et al. (Kandel et al., 2012) Data Analysts Multivariate (categorical and quantitative)
Exploratory data analysis, build statistical models
High: programming expertise
Alspaugh et al. (Alspaugh et al., 2019) Data Analysts Multivariate (categorical and quantitative) Exploratory data analysis High: programming and GUI tools
Demiralp et al. (Demiralp et al., 2017) Data Scientists Multivariate (categorical and quantitative) Exploratory data analysis High: Python programmers
Zhang et al. (Zhang et al., 2022) Business Analysts Multivariate (categorical and quantitative) Making reports and dashboards Medium: limited ML/data science backgrounds
Newburger and Elmquist (Eric Newburger, 2023) Statisticians Multivariate (categorical and quantitative) Exploratory and confirmatory analysis High: scripting to build statistical models
Bartram et al. (Bartram et al., 2021) Data Workers Multivariate (categorical and quantitative) Sensemaking Low: no ML/data science backgrounds
South et al. (South et al., 2020) Political Scientists Text Exploratory analysis and sensemaking Low: no ML/data science backgrounds
Jasim et al. (Jasim et al., 2021) Community Leaders Text Analysis and decision-making Low: no ML/data science backgrounds
Kang and Stasko (Kang and Stasko, 2012) Investigative Analysts Text and documents Sensemaking Medium: limited ML/data science backgrounds
Bhattacharya et al. (Bhattacharya et al., 2024) Healthcare Experts Multivariate (categorical and quantitative) Exploratory model steering Low: no ML/data science backgrounds
Khadpe et al. (Khadpe et al., 2024) Line Managers Nominal/qualitative list, value tree Decision-making Low: no ML/data science backgrounds
Our Work Business Professionals Multivariate (categorical and quantitative) Improve outcomes and decision-making Low; no ML/data science backgrounds
Table 1. Landscape of data, high-level task, and technical engagement of some related research works on data analysis by target audience in a variety of job roles. Our research focuses on business professionals, who are domain experts involved in improving and making decisions using multivariate data but do not engage in technical tasks like programming.
@@ 2.1. Understanding Professionals Performing Data Analysis
Many studies focus on professionals using technical approaches, like data analysts, data scientists, and business analysts, who mainly support decisions rather than directly make them. For example, Kandel et al. (Kandel et al., 2012) identify challenges in analysts’ data pipelines, while Alspaugh et al. (Alspaugh et al., 2019) explore challenges in exploratory analysis. Demiralp et al. (Demiralp et al., 2017) study data scientists’ exploratory analysis, and Zhang et al. investigate business analysts’ needs in generating reports (Zhang et al., 2022) and narratives (Zhang et al., 2020). Newburger and Elmquist study statisticians’ decision-making focusing on visualization (Eric Newburger, 2023). Similar to business professionals, these people work with multivariate data consisting of both categorical and quantitative attributes. However, they have strong technical backgrounds and engage in exploratory or confirmatory data analysis, which is distinct from WIA in decision-making.
@@
Bartram et al. (Bartram et al., 2021) study “data workers” who engage in data analysis as part of their daily tasks but lack formal training in technical domains like computer science, statistics, or machine learning. Previous work often focuses on data workers in non-business domains, such as political science (South et al., 2020), community leadership (Jasim et al., 2021), and law enforcement (Kang and Stasko, 2012). These data workers are similar to business professionals in terms of their low technical expertise. However, many of them work on data exploration and sense-making tasks (Klein et al., 2006; Russell et al., 1993; Klein et al., 2007; Pirolli and Card, 2005) instead of decision-making, and do not discuss processes that involve simulating alternative scenarios and action selection (Lunenburg, 2010; Bruch and Feinberg, 2017; Chen, 2008; Oral et al., 2023). For those data workers who need to make decisions, they primarily work with unstructured data like text and nominal data (Edge et al., 2018; Jasim et al., 2021; Khadpe et al., 2024), which present different challenges from multivariate data. By studying business professionals who perform WIA, our work focuses on a unique combination of the data, task, and technical level dimensions as shown in Table 1.
@@ 2.2. Use of Dashboards in Enterprises
Previous work investigates the role of visualization and dashboards in enterprise data analysis and decision-making. For example, studies find that business professionals, like sales managers, frequently use dashboards (Velcu-Laitinen and Yigitbasioglu, 2012; Bera, 2016; Noonpakdee et al., 2018). However, they are used mainly to communicate findings to stakeholders, not to make data-driven decisions. Dimara and Stasko (Dimara and Stasko, 2021) note the lack of studies on decision-making in visualization research, and other works (Dimara et al., 2021) show how visualizations can be integrated into long-term decisions. Yet, business professionals face challenges in creating dashboards and conducting analysis necessary for making decisions (Sarikaya et al., 2018; Tory et al., 2023). This is due to incomplete information displayed in dashboards and their limited interactivity, which restricts professionals from drilling into the data to explore different scenarios or conduct simulations. Therefore, visualization alone without data-driven analytics is insufficient, leading us to focus on how advanced analytics can support or hinder decision-making.